### v1

# Star clusters in independence complexes of hypergraphs

Abstract.
We study the concept of star clusters in simplicial complexes, which was introduced by Barmak in 2013, by relating it with the structure of hypergraphs that correspond to the simplicial complexes. This implies that for a hypergraph $H$ with a vertex $v$ that is not isolated and not contained in an induced Berge cycle of length $3$, there exists a hypergraph $H^{\prime}$ with fewer vertices than $H$ such that the independence complex of $H$ is homotopy equivalent to the suspension of the independence complex of $H^{\prime}$. As an application, we also prove that if a hypergraph has no Berge cycle of length 00 mod $3$, then the sum of all reduced Betti numbers of its independence complex is at most $1$.
Key words and phrases:
Independence complexes, Hypergraphs, Berge cycles, Star cluster, Homotopy type
2020 Mathematics Subject Classification:
05E45, 55U10
1. Introduction
Throughout the paper, we consider only finite vertex set for graphs, hypergraphs and simplicial complexes. An independent set of a graph $G$ is a vertex subset that does not contain any edge. The independence complex $I(G)$ of $G$ is the collection of all independent sets of $G$. The class of independence complexes of graphs is the same as the class of flag complexes, the simplicial complexes where all minimal non-faces have size $2$. In general, for any finite simplicial complex $X$, there is an independence complex $I(G)$ that is homotopy equivalent to $X$. To see this, consider the barycentric subdivision $\text{sd}(X)$ of $X$. Clearly, $\text{sd}(X)$ is homotopy equivalent to $X$. On the other hand, $\text{sd}(X)$ is a flag complex, and hence it equals to an independence complex of some graph. Thus homotopy types and homology groups of independence complexes of graphs are as various as those of simplicial complexes.
Many results have been obtained on topological properties of independence complexes of graphs, especially on the homotopy types and homology groups of independence complexes of special classes of graphs. For example, Kozlov [14] determined the homotopy types of independence complexes of paths and cycles, and Ehrenborg and Hetyei [9] showed the homotopy types of independence complexes of forests are either contractible or a sphere. In [13], it was proved that the homotopy type of the independence complex of a graph with no induced cycles of length divisible by $3$ are also either contractible or a sphere, strengthening the result by Engström [8]. See also [12] for chordal graphs, [3] for square grids, and [4] for stable Kneser graphs.
Meanwhile, Barmak [1] introduced the notion of star clusters in independence complexes of graphs, which turned out to be useful to understand the independence complexes of graphs. He actually gave alternative proofs of several known results on the independence complexes of graphs by simpler arguments using his star cluster theorem: see Theorem 3.1. We generalize Barmak’s result, providing an analogous statement for the independence complexes of hypergraphs, and then relate it with a hypergraph operation. Using the star cluster theorems, we also establish a hypergraph analogue of a weaker version of the result in [13], showing that the boundedness of the total Betti number of the independence complex of a hypergraph with certain forbidden cycle structure.
1.1. Basic terminology about hypergraphs
A hypergraph $H$ on $V$ is a collection of nonempty subsets of $V$, where each element of $H$ is called an edge of $H$. Note that a graph is a hypergraph whose edges have size exactly $2$. We denote by $V(H)$ the vertex set of $H$. The independent sets and independence complexes of hypergraphs can be defined in analogous way. An independent set of a hypergraph $H$ is a vertex subset that does not contain an edge as a subsets, and the independence complex $I(H)$ of $H$ is a collection of all independent sets of $H$. That is,
$I(H):=\{W\subset V(H)\mid W\text{ contains no edge of }H\}.$
A vertex $v\in V(H)$ is isolated in $H$ if it is not contained in any edge of $H$. Observe that, if $v$ is isolated in $H$, then $I(H)$ is a cone with apex $v$, thus is contractible.
If $H$ has two edges $f$ and $g$ with $g\subset f$, then we clearly have $I(H)=I(H-\{f\})$. In addition, if $\{v\}$ is an edge of $H$ for a vertex $v$, then $I(H)=I(H-v)$ where $H-v$ is the hypergraph obtained from $H$ deleting the vertex $v$ and all edges containing $v$. Therefore, unless stated otherwise, we assume that every edge in a hypergraph is inclusion-minimal and has size at least $2$.
Now we consider “cycles” in hypergraphs. Among various possible definitions of cycles in hypergraphs, we consider “Berge cycle”, which is broadest concept of cycles. A Berge cycle of length $k$ in a hypergraph $H$ is an alternating sequence $C=v_{1}e_{1}v_{2}e_{2}\cdots v_{k}e_{k}$ of $k$ distinct vertices $v_{1},v_{2},\ldots,v_{k}$ and $k$ distinct edges $e_{1},e_{2},\ldots,e_{k}$ of $H$ where $v_{i},v_{i+1}\in e_{i}$ for all index $i$ chosen modulo $k$. Note that in graphs, Berge cycles are equal to usual cycles. We say a Berge cycle $C=v_{1}e_{1}v_{2}e_{2}\cdots v_{k}e_{k}$ in $H$ is induced if $e_{i}\cap\{v_{1},v_{2},\ldots,v_{k}\}=\{v_{i},v_{i+1}\}$ for all $i$ chosen modulo $k$ and there is no edge of $H$ containing two non-consecutive vertices of $C$.
1.2. Star clusters in hypergraphs
Our first main theorem is the following for star clusters in independence complexes of hypergraphs, which generalizes Barmak’s result [1]. Here $\Sigma X$ is the suspension of a simplicial complex $X$. Here, $\text{st}(\sigma)$ of $I(H)$ is the subcomplex whose maximal faces are precisely the maximal faces of $I(H)$ that contain $\sigma$.
Theorem 1.1.
Let $H$ be a hypergraph and $v$ be a vertex of $H$ such that $v$ is not isolated and not contained in an induced Berge cycle of length $3$. Then
$I(H)\simeq\Sigma(\text{st}(v)\cap(\bigcup_{i=1}^{k}\text{st}(\tilde{e}_{i}))),$
where $e_{1},e_{2},\ldots,e_{k}$ are the edges of $H$ containing $v$ and $\tilde{e}_{i}=e_{i}\setminus\{v\}$.
One important remark is that every simplicial complex can be given as an independence complex of some hypergraph. For a simplicial complex $X$ on $V$, let $H_{X}$ be the hypergraph on $V$ where the edges of $H_{X}$ are the minimal non-faces of $X$. Then it is clear that $X$ is equal to the independence complex of $H_{X}$. Thus Theorem 1.1 can be interpreted to a result for arbitrary simplicial complexes.
Also, we can deduce the following from Theorem 1.1.
Theorem 1.2.
Let $H$ be a hypergraph and $v$ be a vertex of $H$ such that $v$ is not isolated and not contained in an induced Berge cycle of length $3$.
  1. (1)
    There is a hypergraph $H^{\prime}$ defined on $V(H)\setminus\{v\}$ such that $I(H)\simeq\Sigma I(H^{\prime})$.
  2. (2)
    $I(H)$ has the homotopy type of a suspension.
  3. (3)
    $\tilde{H}_{1}(I(H))$ is a free abelian group.
1.3. Total Betti numbers of independence complexes of hypergraphs
Our second main theorem is a result on the Betti numbers of independence complexes of certain hypergraphs, which can be obtained as an application of the star cluster theorem. This is inspired by a conjecture by Kalai and Meshulam [11] for graphs, that suggests a connection between the topology of independence complexes and the cycle structure of graphs. They conjectured that for a graph with no induced cycle of length 00 modulo $3$, the number of independent sets of odd size and the number of independent sets of even size differ by at most $1$. Such a graph is said to be ternary. This conjecture was proved in [6] by a purely combinatorial argument.
Indeed, the difference between the number of independent sets of odd size and the number of independent sets of even size in a graph $G$ is actually equal to the reduced Euler characteristic $\tilde{\chi}(I(G))$ of $I(G)$, which is a well-known topological invariant of $I(G)$. In this respect, Kalai and Meshulam also conjectured a stronger version that the sum of all reduced Betti numbers of $I(G)$ is at most $1$ for any ternary graph $G$. This stronger conjecture was proved in [16], and even a stronger statement was proved in [13], arguing that the homotopy type of $I(G)$ of a ternary graph $G$ is either contractible or homotopy equivalent to a sphere.
Going back further than the presentation of the Kalai–Meshulam’s conjectures, the following result by Kozlov [14] on independence complexes of cycles already showed a phenomenon related to modulo $3$. That is, if $C_{\ell}$ is a cycle of length $\ell$, then
$I(C_{\ell})\simeq\begin{cases}\mathbb{S}^{k}\vee\mathbb{S}^{k}&\text{ if }\ell=3k+3,\\ \mathbb{S}^{k}&\text{ if }\ell=3k+2,3k+4,\end{cases}$
where $\mathbb{S}^{k}$ is the $k$-dimensional sphere and $X\vee Y$ is the wedge sum of two topological spaces of $X$ and $Y$. Furthermore, Csorba [7] proved that if $\tilde{G}$ is a graph obtained from $G$ by replacing an edge by a path of length $4$, then $I(\tilde{G})\simeq\Sigma I(G)$, which directly explains the above Kozlov’s result since $\Sigma\mathbb{S}^{k}\simeq\mathbb{S}^{k+1}$. The result by Csorba is actually a special case of the star cluster theorem. Thus it is natural to suspect that the star cluster theorem can be related to the Kalai–Meshulam conjecture.
Considering hypergraphs, we prove the following theorem using the star cluster theorem.
Theorem 1.3.
Let $H$ be a hypergraph with no Berge cycle of length divisible by $3$. Then the sum of all reduced Betti numbers of $I(H)$ is at most $1$.
Organization
In Section 2, we introduce basic terminologies and topological backgrounds. We present the proof of Theorem 1.1 in Section 3 and introduce a specific structural condition for Theorem 1.1 and discuss some examples in Section 4. In Section 5, we prove Theorem 1.3 and also discuss a relation between the cycle structure and bounds on the total Betti number, using the Erdős–Posa theorem for cycles of length 00 mod $3$.
2. Preliminary
In this section, we discuss basic terminologies and topological backgrounds that we need.
A simplicial complex $X$ on a vertex set $V$ is a collection of subsets of $V$ that is closed under taking subsets. Each element of a simplicial complex $X$ is called a face of $X$, and a missing face of $X$ is an inclusion-minimal non-face of $X$.
2.1. Homology groups and Homotopy theory
We introduce basic topological backgrounds on homology groups and homotopy theory. See, for example, [10] for more details.
For a simplicial complex $X$, we denote by $\tilde{H}_{i}(X)$ the $i$-th (reduced) homology group of $X$ over $\mathbb{Z}$. The $i$-th (reduced) Betti number of $X$ is the integer $\tilde{\beta}_{i}(X)=\text{rank}(\tilde{H}_{i}(X))$, and the total Betti number $\beta(X)$ of $X$ is the sum of all reduced Betti numbers of $X$.
We denote by $X\simeq Y$ if two topological spaces $X$ and $Y$ are homotopy equivalent. For a topological space $X$, the suspension $\Sigma X$ of $X$ is the quotient space
$\Sigma X:=X\times[0,1]/(X\times\{0\})/(X\times\{1\}),$
which means $\Sigma X$ is obtained from $X\times[0,1]$ by identifying each of $X\times\{0\}$ and $X\times\{1\}$ as single points. For example, $\Sigma\mathbb{S}^{n}\simeq\mathbb{S}^{n+1}$ for the $n$-dimensional sphere $\mathbb{S}^{n}$.
The following is a well-known fact on the suspension of a simplicial complex.
Proposition 2.1.
For a simplicial complex $X$, $\tilde{H}_{i+1}(\Sigma X)\cong\tilde{H}_{i}(X)$ for all $i$. Hence, $\beta(\Sigma X)=\beta(X)$.
We also need the following known facts in the homotopy theory.
Proposition 2.2.
Let $K,K_{1},K_{2}$ be simplicial complexes such that $K=K_{1}\cup K_{2}$. Then
  1. (a)
    $K/K_{2}\simeq K_{1}/(K_{1}\cap K_{2})$, and
  2. (b)
    for simplicial complexes $M$ and $L$, if $L$ is contractible in $M$, then $M/L\simeq M\vee\Sigma L$. In particular, $M/L\simeq\Sigma L$ if $M$ is contractible, and $M/L\simeq M$ if $L$ is contractible.
By combining (a) and (b), we also obtain the following.
  1. (c)
    Assume both $K_{1}$ and $K_{2}$ are contractible. Then
    $K\simeq K/K_{2}\simeq K_{1}/(K_{1}\cap K_{2})\simeq\Sigma(K_{1}\cap K_{2}).$
    If $K_{1}\cap K_{2}$ is contractible, then $K$ is also contractible.
When we study simplicial complxes, one of the most frequently used tool is the Mayer–Vietoris exact sequence. Let $A,B$ be simplicial complexes and $K=A\cup B$. Then the Mayer–Vietoris sequence for $(K,A,B)$ is the following long exact sequence:
(1) $\cdots\to\tilde{H}_{i}(A\cap B)\to\tilde{H}_{i}(A)\oplus\tilde{H}_{i}(B)\xrightarrow{\gamma_{i}}\tilde{H}_{i}(K)\xrightarrow{\delta_{i}}\tilde{H}_{i-1}(A\cap B)\to\cdots.$
The following inequality about the Betti numbers, which will be used in Section 5, follows from (1).
Lemma 2.3.
For simplicial complexes $A,B$ and $K=A\cup B$, the following holds for all $i$:
$\tilde{\beta}_{i}(K)\leq\tilde{\beta}_{i}(A)+\tilde{\beta}_{i}(B)+\tilde{\beta}_{i-1}(A\cap B).$
Proof.
Consider the Mayer–Vietoris sequence (1) for $(K,A,B)$. Since (1) is exact, we obtain the following short exact sequence for each $i$:
$0\to\text{ker}\delta_{i}\hookrightarrow\tilde{H}_{i}(K)\xrightarrow{\delta_{i}}\text{im}\delta_{i}\to 0$
Then we obtain
$\tilde{\beta}_{i}(K)=\text{rk}(\text{ker}\delta_{i})+\text{rk}(\text{im}\delta_{i})=\text{rk}(\text{im}\gamma_{i})+\text{rk}(\text{im}\delta_{i})\leq\tilde{\beta}_{i}(A)+\tilde{\beta}_{i}(B)+\tilde{\beta}_{i-1}(A\cap B).\qed$
3. Proof of Theorem 1.1 : star cluster theorem
In this section, we prove Theorem 1.1. Before that, we first recall Barmak’s theorem on star clusters in independence complexes of graphs.
For a graph $G$ and a vertex $v$ of $G$, the neighbor $N_{G}(v)$ of $v$ in $G$ is the set of vertices adjacent to $v$ in $G$ and $N_{G}[v]:=N_{G}(v)\cup\{v\}$. For a vertex subset $W$ of $G$, the neighbor $N_{G}(W)$ of $W$ in $G$ is $N_{G}(W):=\bigcup_{w\in W}N_{G}(w)$ and $N_{G}[W]:=N_{G}(W)\cup W$. Let $X$ be a simplicial complex and $\sigma$ be a face of $X$. The star $\text{st}_{X}(\sigma)$ of $\sigma$ in $X$ is defined as
$\text{st}_{X}(\sigma)=\{\tau\subset V\mid\tau\cup\sigma\in X\}.$
Note that $\text{st}_{X}(\sigma)$ is always contractible. The star cluster $\text{SC}_{X}(\sigma)$ of $\sigma$ in $X$ is defined as
$\text{SC}_{X}(\sigma)=\bigcup_{v\in\sigma}\text{st}_{X}(v).$
We write $\text{st}_{X}(\sigma)=\text{st}(\sigma)$ and $\text{SC}_{X}(\sigma)=\text{SC}(\sigma)$ if it is clear which ground complex we are dealing with.
Theorem 3.1 ([1]).
Let $G$ be a graph and $v$ be a vertex that is not isolated and not contained in a triangle. Then
$I(G)\simeq\Sigma(\text{st}(v)\cap SC(N_{G}(v))).$
The above theorem is useful to understand the independence complex of graphs, as $\text{st}(v)\cap SC(N_{G}(v))$ is a simplicial complex defined on the vertex set $V(G)-N_{G}[v]$ that has fewer vertices than $I(G)$. However, it is not guaranteed that $\text{st}(v)\cap SC(N_{G}(v))$ is an independence complex of a graph. For this reason, Theorem 3.1 has limitation in relating the topology of independence complexes and graph structure. On the other hand, $\text{st}(v)\cap SC(N_{G}(v))$ can be viewed as the independence complex of some hypergraph, and this leads us to invent the following hypergraph analogue of Theorem 3.1.
Theorem 1.1.
Let $H$ be a hypergraph and $v$ be a vertex of $H$ such that $v$ is not isolated and not contained in an induced Berge cycle of length $3$. Then
$I(H)\simeq\Sigma(\text{st}(v)\cap(\bigcup_{i=1}^{k}\text{st}(\tilde{e}_{i}))),$
where $e_{1},e_{2},\ldots,e_{k}$ are the edges of $H$ containing $v$ and $\tilde{e}_{i}=e_{i}\setminus\{v\}$.
Note that Theorem 3.1 can be obtained from Theorem 1.1 by replacing $H$ with a graph. In order to prove Theorem 1.1, We first prove the following lemma.
Lemma 3.2.
Let $X$ be a simplicial complex on $V$. Let $\sigma\in X$ and $\sigma_{1},\sigma_{2},\ldots,\sigma_{k}\subset\sigma$. Assume that if $u\in\sigma_{i}\setminus\sigma_{j}$ and $w\in\sigma_{j}\setminus\sigma_{i}$ for some $i,j\in[k]$, then all missing face containing both $u$ and $w$ have a vertex not in $\text{st}_{X}(\sigma_{i})\cup\text{st}_{X}(\sigma_{j})$. Then $\bigcup_{i=1}^{k}\text{st}_{X}(\sigma_{i})$ is contractible.
Proof.
We prove it by the induction on $k$. If $k=1$, then it is trivial. Suppose $k>1$. Let $X_{1}=\bigcup_{i=1}^{k-1}\text{st}_{X}(\sigma_{i})$ and $X_{2}=\text{st}_{X}(\sigma_{k})$. By the induction hypothesis, both $X_{1}$ and $X_{2}$ are contractible. To show $X_{1}\cup X_{2}$ is contractible, it is enough to show $X_{1}\cap X_{2}$ is contractible by Proposition 2.2 (c). Note that $X_{1}\cap X_{2}=\bigcup_{i=1}^{k-1}(\text{st}_{X}(\sigma_{i})\cap\text{st}_{X}(\sigma_{k}))$.
Claim 3.3.
$\text{st}_{X}(\sigma_{i})\cap\text{st}_{X}(\sigma_{k})=\text{st}_{X}(\sigma_{i}\cup\sigma_{k})$ for each $i\in[k-1]$.
Proof of Claim 3.3.
We first show the easy direction that $\text{st}_{X}(\sigma_{i})\cap\text{st}_{X}(\sigma_{k})\supset\text{st}_{X}(\sigma_{i}\cup\sigma_{k})$. If $\eta\in\text{st}_{X}(\sigma_{i}\cup\sigma_{k})$, then $\eta\cup\sigma_{i}\cup\sigma_{k}\in X$. Hence, we obtain $\eta\cup\sigma_{i},\eta\cup\sigma_{k}\in X$, which is equivalent to $\eta\in\text{st}_{X}(\sigma_{i})\cap\text{st}_{X}(\sigma_{k})$.
Now for the opposite direction, take $\tau\in\text{st}_{X}(\sigma_{i})\cap\text{st}_{X}(\sigma_{k})$. We want to show $\tau\in\text{st}_{X}(\sigma_{i}\cup\sigma_{k})$, i.e. $\tau\cup\sigma_{i}\cup\sigma_{k}\in X$. Suppose $\tau\cup\sigma_{i}\cup\sigma_{k}\not\in X$. Then we can take a missing face $S$ of $X$ in $\tau\cup\sigma_{i}\cup\sigma_{k}$. Note that $\tau\cup\sigma_{i},\tau\cup\sigma_{k}\in X$ since $\tau\in\text{st}_{X}(\sigma_{i})\cap\text{st}_{X}(\sigma_{k})$. Then $S\setminus(\tau\cup\sigma_{i})\neq\varnothing$ and $S\setminus(\tau\cup\sigma_{k})\neq\varnothing$. Thus we can take $u\in S\setminus(\tau\cup\sigma_{i})\subset\sigma_{k}\setminus\sigma_{i}$ and $w\in S\setminus(\tau\cup\sigma_{k})\subset\sigma_{i}\setminus\sigma_{k}$. Then by the assumption, $S$ must contain a vertex $x$ not in $\text{st}_{X}(\sigma_{i})\cup\text{st}_{X}(\sigma_{k})$. Since $x\in S\subset\tau\cup\sigma_{i}\cup\sigma_{k}$, we obtain that either $x\in\tau\cup\sigma_{i}$ or $x\in\tau\cup\sigma_{k}$. However, it implies that $x$ is a vertex of $\text{st}_{X}(\sigma_{i})\cup\text{st}_{X}(\sigma_{k})$ since $\tau\cup\sigma_{i},\tau\cup\sigma_{k}\in X$. Thus we reach a contradiction and this completes the proof. ∎
Now, by Claim 3.3, we have $X_{1}\cap X_{2}=\bigcup_{i=1}^{k-1}\text{st}_{X}(\sigma_{i}\cup\sigma_{k})$. To use the induction hypothesis, we will show that $\sigma_{1}\cup\sigma_{k},\sigma_{2}\cup\sigma_{k},\ldots,\sigma_{k-1}\cup\sigma_{k}$ satisfy the assumption in the statement. First, it is obvious that each $\sigma_{i}\cup\sigma_{k}$ is contained in $\sigma$ for $i\in[k-1]$. Furthermore, if $u\in(\sigma_{i}\cup\sigma_{k})\setminus(\sigma_{j}\cup\sigma_{k})\subset\sigma_{i}\setminus\sigma_{j}$ and $w\in(\sigma_{j}\cup\sigma_{k})\setminus(\sigma_{i}\cup\sigma_{k})\subset\sigma_{j}\setminus\sigma_{i}$ for some $i,j\in[k-1]$, then all missing face containing both $u$ and $w$ have a vertex not in $\text{st}_{X}(\sigma_{i})\cup\text{st}_{X}(\sigma_{j})$, and also a vertex not in $\text{st}_{X}(\sigma_{i}\cup\sigma_{k})\cup\text{st}_{X}(\sigma_{j}\cup\sigma_{k})$ since $\text{st}_{X}(\sigma_{i}\cup\sigma_{k})\cup\text{st}_{X}(\sigma_{j}\cup\sigma_{k})\subset\text{st}_{X}(\sigma_{i})\cup\text{st}_{X}(\sigma_{j})$. Thus, by the induction hypothesis, $X_{1}\cap X_{2}$ is contractible. Therefore, Proposition 2.2 (c) implies that $X_{1}\cup X_{2}=\bigcup_{i=1}^{k}\text{st}(\sigma_{i})$ is contractible. ∎
Now we are ready to give a proof of Theorem 1.1.
Proof of Theorem 1.1.
Let $K=\bigcup_{i=1}^{k}\text{st}(\tilde{e}_{i})$. First we show $I(H)=\text{st}(v)\cup K$. We obviously has $\text{st}(v)\cup K\subset I(H)$, and hence it is sufficient show that $I(H)\subset\text{st}(v)\cup K$. Take $\sigma\in I(H)$. If $\sigma\cup\{v\}\in I(H)$, then $\sigma\in\text{st}(v)$. Otherwise, if $\sigma\cup\{v\}\not\in I(H)$, then $\sigma\cup\{v\}$ contains an edge $e$ of $H$. Since $\sigma\in I(H)$, the edge $e$ must contain the vertex $v$. This implies $e=e_{j}$ for some $j\in[k]$, and it follows that $\tilde{e}_{j}\subset\sigma$. Thus $\sigma\cup\tilde{e}_{j}=\sigma\in\text{st}(\tilde{e}_{j})\subset K$. In both cases, we have $I(H)\subset\text{st}(v)\cup K$.
Now, if we show that both $\text{st}(v)$ and $K$ are contractible, then it implies $I(H)\simeq\Sigma(\text{st}(v)\cap K)$ by Proposition 2.2 (c). Clearly $\text{st}(v)$ is contractible, so it is sufficient to show that $K$ is contractible. This follows from an immediate application of Lemma 3.2 to $K=\bigcup_{i=1}^{k}\text{st}(\tilde{e}_{i})$. It remains to check the the following two conditions to apply Lemma 3.2:
  1. (1)
    $\bigcup_{i=1}^{k}\tilde{e}_{i}\in I(H)$.
  2. (2)
    If $u\in\tilde{e}_{i}\setminus\tilde{e}_{j}$ and $w\in\tilde{e}_{j}\setminus\tilde{e}_{i}$, then all missing face containing both $u$ and $w$ have a vertex not in $\text{st}(\tilde{e}_{i})\cup\text{st}(\tilde{e}_{j})$.
To prove (1), assume $\bigcup_{i=1}^{k}\tilde{e}_{i}\not\in I(H)$. Then there is $e\in H$ such that $e\subset\bigcup_{i=1}^{k}\tilde{e}_{i}\subset V(H)\setminus\{v\}$. Recall that every edge in $H$ is inclusion-minimal, so $e\not\subset\tilde{e}_{i}$ for each $i\in[k]$. Now we claim that there is an induced Berge cycle of length $3$ containing $v$, which contradicts the assumption. Since $e\not\subset\tilde{e}_{i}$ for each $i\in[k]$ and $e\subset\bigcup_{i=1}^{k}\tilde{e}_{i}$, we can take two distinct index $i_{1},i_{2}\in[k]$ such that $e\cap(\tilde{e}_{i_{1}}\setminus\tilde{e}_{i_{2}})\neq\varnothing$ and $e\cap(\tilde{e}_{i_{2}}\setminus\tilde{e}_{i_{1}})\neq\varnothing$. Take $u_{1}\in e\cap(\tilde{e}_{i_{1}}\setminus\tilde{e}_{i_{2}})$ and $u_{2}\in e\cap(\tilde{e}_{i_{2}}\setminus\tilde{e}_{i_{1}})$. Note that $e\cap\{v,u_{1},u_{2}\}=\{u_{1},u_{2}\}$, ${e}_{i_{1}}\cap\{v,u_{1},u_{2}\}=\{v,u_{1}\}$, and ${e}_{i_{2}}\cap\{v,u_{1},u_{2}\}=\{v,u_{2}\}$. Hence, $v~{}e_{i_{1}}~{}u_{1}~{}e~{}u_{2}~{}e_{i_{2}}$ is an induced Berge cycle of length $3$ containing $v$. Therefore, we have $\bigcup_{i=1}^{k}\tilde{e}_{i}\not\in I(H)$.
Now we prove (2). First, note that the missing faces of $I(H)$ are exactly the edges of $H$. Suppose the contrary that there is a missing face $e$ of $I(H)$, i.e. an edge $e$ of $H$, containing both $u$ and $w$ such that all vertices in $e$ are contained in $\text{st}(\tilde{e}_{i})\cup\text{st}(\tilde{e}_{j})$. Note that $v$ is not a vertex of $\text{st}(\tilde{e}_{i})\cup\text{st}(\tilde{e}_{j})$, hence we have $v\not\in e$. Similarly as in the proof of (1), $v~{}e_{i}~{}u~{}e~{}w~{}e_{j}$ is an induced Berge cycle of length $3$ containing $v$. This is a contradiction, thus shows (2). ∎
Now we recall Theorem 1.2, which can be obtained from Theorem 1.1.
Theorem 1.2.
Let $H$ be a hypergraph and $v$ be a vertex of $H$ such that $v$ is not isolated and not contained in an induced Berge cycle of length $3$.
  1. (1)
    There is a hypergraph $H^{\prime}$ with $V(H^{\prime})\subset V(H)\setminus\{v\}$ such that $I(H)\simeq\Sigma I(H^{\prime})$.
  2. (2)
    $I(H)$ has the homotopy type of a suspension.
  3. (3)
    $\tilde{H}_{1}(I(H))$ is a free abelian group.
Proof.
Let $e_{1},e_{2},\ldots,e_{k}$ be the edges of $H$ containing $v$ and $\tilde{e}_{i}=e_{i}\setminus\{v\}$. By Theorem 1.1, we know $I(H)\simeq\Sigma(\text{st}(v)\cap(\bigcup_{i=1}^{k}\text{st}(\tilde{e}_{i})))$. Let $K=\text{st}(v)\cap(\bigcup_{i=1}^{k}\text{st}(\tilde{e}_{i}))$. To prove (1), it is enough to see that $K$ is a simplicial complex defined on a subset of $V(H)\setminus\{v\}$. This is obvious since each $\text{st}(\tilde{e}_{i})$ is a complex on $V(H)\setminus\{v\}$. For (2), it immediately follows from (1). For (3), since $I(H)\simeq\Sigma K$, we have $\tilde{H}_{1}(I(H))\cong\tilde{H}_{0}(K)$ by Proposition 2.1. $\tilde{H}_{0}(K)$ is always a free abelian group, thus it implies that $\tilde{H}_{1}(I(H))$ is also a free abelian group. ∎
4. Structural Condition for the Star cluster Theorem
Let $H$ be a hypergraph on $V$ and $v$ be a vertex of $H$ that is not isolated and not contained in an induced Berge cycle of length $3$. Let $e_{1},e_{2},\ldots,e_{k}$ be the edges of $H$ containing $v$ and $\tilde{e}_{i}=e_{i}\setminus\{v\}$. Define $H_{v}$ be the set of missing faces of $st(v)\cap(\bigcup_{i=1}^{k}\text{st}(\tilde{e}_{i}))$. Then, by Theorem 1.1, we have $I(H)\simeq\Sigma I(H_{v})$. In this section, we provide a characterization of the edges of $H_{v}$.
Theorem 4.1.
Let $H$ and $H_{v}$ be defined as above. Then we have $\tilde{e}_{1},\tilde{e}_{2},\ldots,\tilde{e}_{k}\in H_{v}$. Furthermore, $f\in H_{v}\setminus\{\tilde{e}_{1},\tilde{e}_{2},\ldots,\tilde{e}_{k}\}$ if and only if $\tilde{e}_{i}\not\subset f$ for all $i\in[k]$ and the following holds.
  1. (i)
    For each $i\in[k]$, there is an edge $h_{i}\in H\setminus\{e_{1},e_{2},\ldots,e_{k}\}$ such that $h_{i}\subset f\cup e_{i}$, and
  2. (ii)
    $f$ is inclusion-minimal subject to (i).
The proof will be given in Section 4.1. Using the above characterization of the edges of $H_{v}$, we describe how to construct $H_{v}$ from the original hypergraph $H$.
Corollary 4.2.
$H_{v}$ can be obtained by the following process.
  1. (1)
    Remove $v$ from each $e_{i}$, that is, instead of $e_{i}$ add $\tilde{e}_{i}=e_{i}\setminus\{v\}$ as a new element.
  2. (2)
    All edges not containing $v$ remain the same.
  3. (3)
    Take $f_{i}\in H\setminus\{e_{1},e_{2},\ldots,e_{k}\}$ such that $f_{i}\cap e_{i}\neq\varnothing$ for each $i\in[k]$. Let $f=\bigcup_{i=1}^{k}(f_{i}\setminus e_{i})$. We add $f$ as a new element.
  4. (4)
    Let $H_{v}$ be the set of all inclusion-minimal elements among all elements obtained by (1)–(3).
For instance, in Figure 1, we obtain more than $1$ new elements at the step (3) and we may ignore singleton edges. In Figure 2, we apply the above process for a $3$-uniform hypergraph.
Figure 1. By applying Theorem 4.1 to $H$ and $v$, we obtain that $H_{v}$ is isomorphic to $K_{2,2}$.
Figure 2. We obtain $H_{v}$ by applying Theorem 4.1 for a $3$-uniform hypergraph $H$ and $v$.
Proof of Corollary 4.2.
We first need to check that all edges of $H_{v}$ appear during the process (1)–(3). Obviously, $\tilde{e}_{1},\ldots,\tilde{e}_{k}\in H_{v}$ appear in the step (1). By Theorem 4.1, all other edges in $f\in H_{v}\setminus\{\tilde{e}_{1},\ldots,\tilde{e}_{k}\}$ satisfies that
  1. (i)
    For each $i\in[k]$, there is an edge $h_{i}\in H\setminus\{e_{1},e_{2},\ldots,e_{k}\}$ such that $h_{i}\subset f\cup e_{i}$, and
  2. (ii)
    $f$ is inclusion-minimal subject to (i).
Suppose $h_{i}\cap e_{i}=\varnothing$ for some $i\in[k]$. Then $h_{i}\subset f\cup e_{i}$ implies $h_{i}\subset f$. In addition, since $h_{i}\subset h_{i}\cup e_{j}$ for all $j\in[k]$, we have $h_{i}=f$ by the minimality of $f$. Thus, in this case, we obtain $f=h_{i}\in H\setminus\{e_{1},e_{2},\ldots,e_{k}\}$. Hence, $f$ appears in the step (2).
Thus assume $h_{i}\cap e_{i}\neq\varnothing$ for all $i\in[k]$. Then define $h=\bigcup_{i=1}^{k}(h_{i}\setminus e_{i})$. For each $i\in[k]$, we have $h_{i}\setminus e_{i}\subset f$ since $h_{i}\subset f\cup e_{i}$, hence it implies $h\subset f$. On the other hand, since $h_{i}\subset h\cup e_{i}$ for all $i\in[k]$, we obtain $h=f$ by the minimality of $f$. Thus, $f=h=\bigcup_{i=1}^{k}(h_{i}\setminus e_{i})$ appears in the step (3).
Now it remains to check that all elements appearing during the process (1)–(3) contain some edge of $H_{v}$. Let $e$ be an element appearing during the process (1)–(3).
  • If $e$ appears in the step (1), then $e=\tilde{e}_{i}$ for some $i$ and we know $e=\tilde{e}_{i}\in H_{v}$ by Theorem 4.1.
  • If $e$ appears in the step (2), then $e\in H\setminus\{e_{1},\ldots,e_{k}\}$. In this case, we have $\tilde{e}_{i}\not\subset e$ and $e$ satisfies the condition (i) of Theorem 4.1 since $e\subset e\cup e_{i}$ for all $i\in[k]$. Thus by Theorem 4.1, $e$ contains some edge of $H_{v}$.
  • If $e$ appears in the step (3), then $e=\bigcup_{i=1}^{k}(f_{i}\setminus e_{i})$ for some $f_{i}\in H\setminus\{e_{1},\ldots,e_{k}\}$ such that $f_{i}\cap e_{i}\neq\varnothing$. Then $e$ satisfies the condition (i) of Theorem 4.1 since $f_{i}\subset e\cup e_{i}$ for each $i\in[k]$, and hence $e$ contains some edge of $H_{v}$ by Theorem 4.1.
This completes the proof. ∎
Based on Corollary 4.2, Theorem 1.1 can be applied to compute the homotopy type of the independence complex of hypergraphs. Here, we introduce one interesting application of Theorem 1.1.
Let $P_{n,k}$ be a $k$-uniform tight path on $n$ vertices, that is, the edges of $P_{n,k}$ are the set of $k$ consecutive vertices of the vertices $v_{1},v_{2},\ldots,v_{n}$ of $P_{n,k}$. Let $H$ be a hypergraph containing $P_{2k-2,k}$ defined on $a_{1},a_{2},\ldots,a_{2k-2}$ as a subgraph. Let $\mathcal{L}_{k}(H)$ be the hypergraph obtained from $H$ by deleting all edges of $P_{2k-2,k}$ on $a_{1},a_{2},\ldots,a_{2k-2}$ and adding new vertices $b_{1},b_{2},\ldots,b_{k+1}$ and new edges of $P_{3k-1,k}$ on $a_{1},\ldots,a_{k-1},b_{1},\ldots,b_{k+1},a_{k},\ldots,a_{2k-2}$. Note that $H$ can have edges on $a_{1},a_{2},\ldots,a_{2k-2}$ incomparable with any edge of $P_{2k-2,k}$. For instance, see Figure 3 when $k=3$.
Figure 3. For a $3$-uniform tight path $P_{6,3}$, we obtain $\mathcal{L}_{3}(P_{6,3})=P_{10,3}$.
If $H$ is a graph, then $k=2$ and $P_{2k-2,k}=P_{2,2}$ is a path of length $1$ on $a_{1},a_{2}$ and $\mathcal{L}_{2}(H)$ is obtained from $H$ by replacing an edge $a_{1}a_{2}$ by a path of length $4$. As mentioned in the introduction, Csorba [7] showed that $I(\mathcal{L}_{2}(H))\simeq\Sigma I(H)$. The following theorem generalizes Csorba’s result.
Theorem 4.3.
Let $H$ and $\mathcal{L}_{k}(H)$ be defined as above. Then $I(\mathcal{L}_{k}(H))\simeq\Sigma^{k-1}I(H)$.
Proof.
We prove by applying the star cluster theorem to $\mathcal{L}_{k}(H)$ for vertices $b_{k},b_{k-1},\ldots,b_{2}$ in order. Let $F_{1}=\mathcal{L}_{k}(H)_{b_{k}}$ and $F_{i}=(F_{i-1})_{b_{k-i+1}}$ for $12\leq i\leq k-$ 1. Then $F_{i}$ is the hypergraph obtained by applying the star cluster theorem to $\mathcal{L}_{k}(H)$ for vertices $b_{k},b_{k-1},\ldots,b_{k-i+1}$ in order.
Claim 4.4.
For each $i\in[k-1]$, the edges of $F_{i}$ induced by $a_{1},a_{2},\ldots,a_{2k-2},b_{1},b_{2},\ldots,b_{k-i},b_{k+1}$ are given as follows:
  1. a)
    All edges of $P_{2k-2i,k-i}$ on $b_{1},b_{2},\ldots,b_{k-i},b_{k+1},a_{k},\ldots,a_{2k-i-2}$ are edges of $F_{i}$.
  2. b)
    $\{a_{j},a_{j+1}\ldots,a_{k-1},b_{1},b_{2},\ldots,b_{j}\}\in F_{i}$ for $11\leq j\leq k-i-$ 1.
  3. c)
    $\{a_{\ell},\ldots,a_{k-1},b_{1},\ldots,b_{m},a_{k+m},\ldots,a_{k+\ell-1}\}\in F_{i}$ for $k-i\leq\ell\leq k-1$, $10\leq m\leq k-i-$ 1. Note that we obtain $\{a_{\ell},a_{\ell+1},\ldots,a_{k+\ell-1}\}$ when $m=0$.
  4. d)
    All other edges on $a_{1},a_{2},\ldots,a_{2k-2}$ in $H$ incomparable with any edge of $P_{2k-2,k}$.
Proof of Claim 4.4.
We prove by the induction on $i$ using Theorem 4.1 and Corollary 4.2. For $i=1$, the edges of $F_{1}$ induced by $a_{1},a_{2},\ldots,a_{2k-2},b_{1},b_{2},\ldots,b_{k-1},b_{k+1}$ are given as follows: We first consider the edges $e_{1},\ldots,e_{k}$ of $\mathcal{L}_{k}(H)$ containing $b_{k}$. Let $e_{1}=\{b_{1},\ldots,b_{k}\}$, $e_{2}=\{b_{2},\ldots,b_{k+1}\}$, and $e_{j}=\{b_{j},\ldots,b_{k+1},a_{k},\ldots,a_{k+j-3}\}$ for $3\leq j\leq k$.
  • In the step (1), we obtain $\tilde{e}_{1},\ldots,\tilde{e}_{k}$, i.e. the edges of $P_{2k-2,k-1}$ on $b_{1},\ldots,b_{k-1},b_{k+1},a_{k},\ldots,a_{2k-3}$.
  • In the step (2), we obtain $\{a_{j},\ldots,a_{k-1},b_{1},\ldots,b_{j}\}$ for $11\leq j\leq k-$ 1 and all other edges on $a_{1},a_{2},\ldots,a_{2k-2}$ in $H$ incomparable with any edge of $P_{2k-2,k}$.
  • In the step (3), we obtain $\{a_{k-1},b_{1},\ldots,b_{m},a_{k+m},\ldots,a_{2k-2}\}$ for $20\leq m\leq k-$ 2 by taking $f_{j}=\{a_{k-1},b_{1},\ldots,b_{k}\}$ for $11\leq j\leq m+$ 1 and $f_{j}=\{b_{k+1},a_{k},\ldots,a_{2k-2}\}$ for $m+2\leq j\leq k$.
Note that $\{a_{k-1},b_{1},\ldots,b_{k-1}\}$ is not minimal since we obtained $\{b_{1},\ldots,b_{k-1}\}$ in the step (1), and all other elements obtained from (1)–(3) is minimal, hence we obtain that the statement holds when $i=1$.
Now suppose $i>1$ and $F_{i-1}$ satisfies the statement, that is, the edges of $F_{i-1}$ induced by
$a_{1},a_{2},\ldots,a_{2k-2},b_{1},b_{2},\ldots,b_{k-i+1},b_{k+1}$ are given as follows:
  1. a)
    All edges of $P_{2k-2i+2,k-i+1}$ on $b_{1},b_{2},\ldots,b_{k-i+1},b_{k+1},a_{k},\ldots,a_{2k-i-1}$ are edges of $F_{i-1}$.
  2. b)
    $\{a_{j},a_{j+1}\ldots,a_{k-1},b_{1},b_{2},\ldots,b_{j}\}\in F_{i-1}$ for $1\leq j\leq k-i$.
  3. c)
    $\{a_{\ell},\ldots,a_{k-1},b_{1},\ldots,b_{m},a_{k+m},\ldots,a_{k+\ell-1}\}\in F_{i-1}$ for $k-i+1\leq\ell\leq k-1$, $0\leq m\leq k-i$.
  4. d)
    All other edges on $a_{1},a_{2},\ldots,a_{2k-2}$ in $H$ incomparable with any edge of $P_{2k-2,k}$.
By Theorem 4.1 and Corollary 4.2, we obtain the edges of $F_{i}=(F_{i-1})_{b_{k-i+1}}$ induced by $a_{1},a_{2},\ldots,a_{2k-2}$, $b_{1},b_{2},\ldots,b_{k-i},b_{k+1}$ are given as follows: Consider the edges $e_{1},\ldots,e_{k-i+1}$ of $F_{i-1}$ containing $b_{k-i+1}$. Let $e_{1}=\{b_{1},\ldots,b_{k-i+1}\}$, $e_{2}=\{b_{2},\ldots,b_{k-i+1},b_{k+1}\}$ and $e_{j}=\{b_{j},\ldots,b_{k-i+1},b_{k+1},a_{k},\ldots,a_{k+j-3}\}$ for $13\leq j\leq k-i+$ 1.
  • In the step (1), we obtain $\tilde{e}_{1},\ldots,\tilde{e}_{k-i+1}$, i.e. the edges of $P_{2k-2i,k-i}$ on $b_{1},b_{2},\ldots,b_{k-i},b_{k+1}$, $a_{k},\ldots,a_{2k-i-2}$.
  • In the step (2), we obtain $\{a_{j},\ldots,a_{k-1},b_{1},\ldots,b_{j}\}$ for $1\leq j\leq k-i$,
    $\{a_{\ell},\ldots,a_{k-1},b_{1},\ldots,b_{m},a_{k+m},\ldots,a_{k+\ell-1}\}$ for $k-i+1\leq\ell\leq k-1$, $0\leq m\leq k-i$, and
    all other edges on $a_{1},a_{2},\ldots,a_{2k-2}$ in $H$ incomparable with any edge of $P_{2k-2,k}$.
  • In the step (3), we obtain $\{a_{k-i},\ldots,a_{k-1},b_{1},\ldots,b_{m},a_{k+m},\ldots,a_{2k-i-1}\}$ for $10\leq m\leq k-i-$ 1 by taking $f_{j}=\{a_{k-i},\ldots,a_{k-1},b_{1},\ldots,b_{k-i}\}$ for $11\leq j\leq m+$ 1 and $f_{j}=\{b_{k+1},a_{k},\ldots,a_{2k-i-1}\}$ for $m+2\leq j\leq k-i+1$.
Note that $\{a_{\ell},\ldots,a_{k-1},b_{1},\ldots,b_{k-i},a_{2k-i},\ldots,a_{k+\ell-1}\}$ for each $k-i\leq\ell\leq k-1$ is not minimal since we obtained $\{b_{1},\ldots,b_{k-i}\}$ in the step (1), and all other elements obtained from (1)–(3) is minimal. Thus we can conclude that $F_{i}$ also satisfies the statement. ∎
By Claim 4.4, we obtain that the edges of $F_{k-1}$ on $a_{1},a_{2},\ldots,a_{2k-2},b_{1},b_{k+1}$ is given as follows:
  1. a)
    All edges of $P_{2,1}$ on $b_{1},b_{k+1}$ are edges of $F_{k-1}$, that is, $\{b_{1}\},\{b_{k+1}\}\in F_{k-1}$.
  2. c)
    $\{a_{\ell},a_{\ell+1},\ldots,a_{k+\ell-1}\}\in F_{k-1}$ for $11\leq\ell\leq k-$ 1.
  3. d)
    All other edges on $a_{1},a_{2},\ldots,a_{2k-2}$ in $H$ incomparable with any edge of $P_{2k-2,k}$.
Thus we obtain that $F_{k-1}=H\cup\{\{b_{1}\},\{b_{k+1}\}\}$. Since $b_{1},b_{k+1}$ are forming singleton edges, they are not appear in $I(F_{k-1})$. Hence, we can conclude that $I(F_{k-1})=I(H)$. Now, by Theorem 1.1, we obtain
$I(\mathcal{L}_{k})\simeq\Sigma I(F_{1})\simeq\Sigma^{k-1}F_{k-1}\simeq\Sigma^{k-1}I(H)$
as we wanted. ∎
4.1. Proof of Theorem 4.1
For each $i\in[k]$, define the hypergraph $H_{i}$ on $V\setminus\{v\}$ as
$H_{i}=\{e\setminus e_{i}:e\in H\setminus\{e_{1},e_{2},\ldots,e_{k}\}\}\cup\{\tilde{e}_{1},\tilde{e}_{2},\ldots,\tilde{e}_{k}\}.$
Note that edges in $H_{i}$ may not be inclusion-minimal.
Lemma 4.5.
Let $H$ and $H_{i}$ be defined as above. Then $I(H_{i})=\text{st}(v)\cap\text{st}(\tilde{e}_{i})$.
Proof.
First we show $I(H_{i})\subset\text{st}(v)\cap\text{st}(\tilde{e}_{i})$. Take $\sigma\in I(H_{i})$. Suppose $\sigma\cup\{v\}\not\in I(H)$. Then there is $e\in H$ such that $e\subset\sigma\cup\{v\}$. If $e\neq e_{j}$ for all $j\in[k]$, then $e\setminus e_{i}\subset(\sigma\cup\{v\})\setminus e_{i}\subset\sigma$. Since $e\setminus e_{i}$ is an edge of $H_{i}$, this is a contradiction to $\sigma\in I(H_{i})$. If $e=e_{j}$ for some $j\in[k]$, then we have $\tilde{e}_{j}\subset\sigma$. Similarly as above, we again obtain an edge $\tilde{e}_{j}$ of $H_{i}$ contained in $\sigma$, which is a contradiction to $\sigma\in I(H_{i})$. This implies $\sigma\cup\{v\}\in I(H)$, that is, $\sigma\in\text{st}(v)$. Next, suppose $\sigma\cup\tilde{e}_{i}\not\in I(H)$. Then there is $e\in H$ such that $e\subset\sigma\cup\tilde{e}_{i}$. Note that $e\neq e_{j}$ for all $j\in[k]$ since $\sigma,\tilde{e}_{i}\subset V\setminus\{v\}$. Thus we have $e\setminus e_{i}\subset(\sigma\cup\tilde{e}_{i})\setminus e_{i}\subset\sigma$. Since $e\setminus e_{i}\in H_{i}$, it is a contradiction to $\sigma\in I(H_{i})$, implying $\sigma\in\text{st}(\tilde{e}_{i})$. Therefore, we have $I(H_{i})\subset\text{st}(v)\cap\text{st}(\tilde{e}_{i})$.
Now we prove $I(H_{i})\supset\text{st}(v)\cap\text{st}(\tilde{e}_{i})$. Take $\sigma\in\text{st}(v)\cap\text{st}(\tilde{e}_{i})$. First, note that $\sigma$ does not contain $v$ since $v\notin\text{st}(\tilde{e}_{i})$. Also, since $\sigma\in\text{st}(v)$ and no $\tilde{e}_{j}$ is contained in $\text{st}(v)$, we have $\tilde{e}_{j}\not\subset\sigma$ for all $j\in[k]$. Now suppose $\sigma\notin I(H_{i})$. Then there is $e\in H_{i}$ such that $e\subset\sigma$. Observing $e\neq\tilde{e}_{j}$ for all $j\in[k]$, it must be $e=f\setminus e_{i}$ for some $f\in H\setminus\{e_{1},e_{2},\ldots,e_{k}\}$. Then $f\subset e\cup\tilde{e}_{i}\subset\sigma\cup\tilde{e}_{i}$, and this is a contradiction to $\sigma\in\text{st}(\tilde{e}_{i})$. Thus, we have $\sigma\in I(H_{i})$ and hence we conclude that $I(H_{i})\supset\text{st}(v)\cap\text{st}(\tilde{e}_{i})$. This completes the proof. ∎
For a hypergraph $F$, let $\text{cl}(F)$ be the closure of $F$ and define it as
$\text{cl}(F)=\{e\subset V:f\subset e\text{ for some }f\in F\}.$
Then we have $\text{cl}(F)=2^{V}\setminus I(F)$ for any hypergraph $F$. If all edges of $F$ are inclusion-minimal, then $f\in F$ if and only if $f$ is inclusion-minimal in $\text{cl}(F)$.
Lemma 4.6.
Let $F,F_{1},F_{2},\ldots,F_{m}$ be hypergraphs on $V$ and $I(F)=\bigcup_{i=1}^{m}I(F_{i})$. Assume that all edges in $F$ are inclusion-minimal. Then $f\in F$ if and only if the following holds.
  1. (1)
    There is an edge $e_{i}\subset f$ of $F_{i}$ for each $i\in[m]$, and
  2. (2)
    $f$ is inclusion-minimal subject to (1).
Proof.
First, we know that $I(F)=\bigcup_{i=1}^{m}I(F_{i})$ implies
$\text{cl}(F)=2^{V}\setminus I(F)=2^{V}\setminus\bigcup_{i=1}^{m}I(F_{i})=\bigcap_{i=1}^{m}2^{V}\setminus I(F_{i})=\bigcap_{i=1}^{m}\text{cl}(F_{i}).$
Thus, we obtain that $f\in F$ if and only if $f$ is inclusion-minimal in $\bigcap_{i=1}^{m}\text{cl}(F_{i})$. Thus it is enough to show that every element $f$ in $\bigcap_{i=1}^{m}\text{cl}(F_{i})$ satisfies the condition that there is an edge $e_{i}\subset f$ of $F_{i}$ for each $i\in[m]$, and this directly follows from the definition of $\text{cl}(F_{i})$. ∎
Now we prove Theorem 4.1.
Theorem 4.1.
We have $\tilde{e}_{1},\tilde{e}_{2},\ldots,\tilde{e}_{k}\in H_{v}$. Furthermore, $f\in H_{v}\setminus\{\tilde{e}_{1},\tilde{e}_{2},\ldots,\tilde{e}_{k}\}$ if and only if $\tilde{e}_{i}\not\subset f$ for all $i\in[k]$ and the following holds.
  1. (i)
    For each $i\in[k]$, there is an edge $h_{i}\in H\setminus\{e_{1},e_{2},\ldots,e_{k}\}$ such that $h_{i}\subset f\cup e_{i}$, and
  2. (ii)
    $f$ is inclusion-minimal subject to (i).
Proof.
Let $H_{i}$ be defined as above. Note that $I(H_{v})=\bigcup_{i=1}^{k}I(H_{i})$ by Theorem 1.1 and Lemma 4.5. We will prove the statement by applying Lemma 4.6 to $I(H_{v})=\bigcup_{i=1}^{k}I(H_{i})$. By Lemma 4.6, $f\in H_{v}$ if and only if
  1. (1)
    there is an edge $f_{i}\subset f$ of $H_{i}$ for each $i\in[k]$, and
  2. (2)
    $f$ is inclusion-minimal subject to (1).
We first prove that $\tilde{e}_{j}\in H_{v}$ for each $j\in[k]$. Since $\tilde{e}_{j}$ itself is an edge of $H_{i}$, thus it contains an edge of $H_{i}$ for all $i\in[k]$. To show the minimality of $\tilde{e}_{j}$, suppose there is a proper subset $g$ of $\tilde{e}_{j}$ such that $g$ contains an edge $h_{j}$ of $H_{j}$. If $h_{j}=\tilde{e}_{i}$ for some $i\in[k]$, then we have $\tilde{e}_{i}=h_{j}\subset g\subsetneq\tilde{e}_{j}$, and hence we obtain $e_{i}\subsetneq e_{j}$ and it contradicts the assumption that every edge in $H$ is inclusion-minimal. Thus $h_{j}=e\setminus e_{j}$ for some $e\in H\setminus\{e_{1},e_{2},\ldots,e_{k}\}$. Then we obtain $e\setminus e_{j}\subset g\subsetneq\tilde{e}_{j}$, and it implies that $e\subsetneq e_{j}$, which is again a contradiction to the assumption that every edge in $H$ is inclusion-minimal. Thus we obtain that $\tilde{e}_{1},\tilde{e}_{2},\ldots,\tilde{e}_{k}\in H_{v}$.
Now to prove the remaining part, we claim the following.
Claim 4.7.
Suppose $f\subset V(H)\setminus\{v\}$ and $\tilde{e}_{j}\not\subset f$ for all $j\in[k]$. Then there is an edge $f_{i}\subset f$ of $H_{i}$ if and only if there is an edge $h_{i}\in H\setminus\{e_{1},e_{2},\ldots,e_{k}\}$ such that $h_{i}\subset f\cup e_{i}$.
Proof.
Suppose that there is an edge $f_{i}\subset f$ of $H_{i}$. Since $\tilde{e}_{j}\not\subset f$ for all $j\in[k]$, we know $f_{i}\neq\tilde{e}_{j}$ for all $j\in[k]$. Thus there is an edge $h_{i}\in H\setminus\{e_{1},e_{2},\ldots,e_{k}\}$ such that $f_{i}=h_{i}\setminus e_{i}$, and hence $h_{i}\subset f\cup e_{i}$. Now for the opposite, if there is an edge $h_{i}\in H\setminus\{e_{1},e_{2},\ldots,e_{k}\}$ such that $h_{i}\subset f\cup e_{i}$, then obviously $h_{i}\setminus e_{i}\in H_{i}$ and $h_{i}\setminus e_{i}\subset f\setminus e_{i}\subset f$. ∎
To prove the only if part, assume that $f\in H_{v}\setminus\{\tilde{e}_{1},\tilde{e}_{2},\ldots,\tilde{e}_{k}\}$. Then by the minimality of each edge of $H_{v}$, we have $\tilde{e}_{i}\not\subset f$ for all $i\in[k]$. In addition, by Lemma 4.6 and Claim 4.7, it is obvious that $f$ satisfies (i),(ii).
Now, to prove the if part, assume that $\tilde{e}_{i}\not\subset f$ for all $i\in[k]$ and $f$ satisfies (i),(ii). Then it is obvious that $f\not\in\{\tilde{e}_{1},\tilde{e}_{2},\ldots,\tilde{e}_{k}\}$. In addition, by Claim 4.7, the condition (i) is equivalent to that there is an edge $f_{i}\subset f$ of $H_{i}$ for each $i\in[k]$. Then by Lemma 4.6, we obtain $f\in H_{v}$, and hence we have $f\in H\setminus\{\tilde{e}_{1},\tilde{e}_{2},\ldots,\tilde{e}_{k}\}$. This completes the proof. ∎
5. Bounds on the total Betti numbers
In this section, we present the proof of Theorem 1.3. Moreover, we show that if a hypergraph $H$ has only bounded number of “disjoint” ternary Berge cycles, then $b(I(H))$ is also bounded. The definition of “disjoint” Berge cycles is given as follows:
We say two Berge cycles $C_{1}=v_{1}f_{1}v_{2}f_{2}\ldots v_{k}f_{k}$, $C_{2}=w_{1}g_{1}w_{2}g_{2}\ldots w_{\ell}g_{\ell}$ are disjoint if $\{v_{1},v_{2},\ldots,v_{k}\}\cap\{w_{1},w_{2},\ldots,w_{\ell}\}=\varnothing$ and $\{f_{1},f_{2},\ldots,f_{k}\}\cap\{g_{1},g_{2},\ldots,g_{\ell}\}=\varnothing$. For a hypergraph $H$, let $t(H)$ be the maximum integer $k$ such that $H$ has $k$ pairwise disjoint ternary Berge cycles. Note that $t(H)=0$ if and only if $H$ has no ternary Berge cycle.
5.1. Proof of Theorem 1.3
Let $H$ be a hypergraph and $e=\{v_{1},v_{2},\ldots,v_{k}\}$ be an edge of $H$. We define a new hypergraph $H_{e}$ obtained from $H$ by deleting the edge $e$ and adding new vertices $\{w,u_{1},u_{2},\ldots,u_{k}\}$ and new edges $\{\{w,u_{1}\},\{w,u_{2}\},\ldots,\{w,u_{k}\},\{u_{1},v_{1}\},\{u_{2},v_{2}\},\ldots,\{u_{k},v_{k}\}\}$. That is,
$H_{e}=(H\setminus\{e\})\cup\{\{w,u_{1}\},\{w,u_{2}\},\ldots,\{w,u_{k}\},\{u_{1},v_{1}\},\{u_{2},v_{2}\},\ldots,\{u_{k},v_{k}\}\}.$
See Figure 4 for an example with $k=4$.
Figure 4. The hypergraph $H_{e}$ for $e=\{v_{1},v_{2},v_{3},v_{4}\}$.
Lemma 5.1.
Let $H$ be a hypergraph and $e$ be an edge of $H$. Then $I(H_{e})\simeq\Sigma I(H)$.
Proof.
We claim that $(H_{e})_{w}=H\cup\{\{u_{1}\},\{u_{2}\},\ldots,\{u_{k}\}\}$. Note that $I((H_{e})_{w})=I(H)$ since all vertices in the singleton edges cannot appear in the independence complex. Then we can conclude that $I(H_{e})\simeq\Sigma I((H_{e})_{w})=\Sigma I(H)$ by Theorem 1.1.
Thus it is enough to show $(H_{e})_{w}=H\cup\{\{u_{1}\},\{u_{2}\},\ldots,\{u_{k}\}\}$, and for that, we will use Corollary 4.2. By Corollary 4.2, we obtain the edges of $(H_{e})_{w}$ as follows:
  • In the step (1), we obtain $\{u_{1}\},\{u_{2}\},\ldots,\{u_{k}\}$ since $\{w,u_{1}\},\{w,u_{2}\},\ldots,\{w,u_{k}\}$ are the edges containing $w$ in $H_{e}$.
  • In the step (2), we obtain all edges of $H_{e}$ not containing $w$, that means we obtain all edges of $H$ except $e$ and $\{u_{1},v_{1}\},\{u_{2},v_{2}\},\ldots,\{u_{k},v_{k}\}$.
  • In the step (3), since $\{u_{i},v_{i}\}$ is the only edge in $H_{e}\setminus\{\{w,u_{1}\},\{w,u_{2}\},\ldots,\{w,u_{k}\}\}$ intersects with $\{w,u_{i}\}$, we only obtain $e=\bigcup_{i=1}^{k}\{v_{i}\}$.
Now we collect only inclusion-minimal elements among obtained by (1)–(3), then we obtain $(H_{e})_{w}=H\cup\{\{u_{1}\},\{u_{2}\},\ldots,\{u_{k}\}\}$ as we wanted. ∎
Lemma 5.2.
For a positive integer $k$, if $t(H)<k$, then $t(H_{e})<k$.
Proof.
Suppose $t(H_{e})\geq k$, that is, $H_{e}$ has $k$ pairwise disjoint ternary Berge cycles $C_{1},C_{2},\ldots,C_{k}$. If all edges in $C_{1},C_{2},\ldots,C_{k}$ are edges of $H$, then $t(H)\geq k$ which is a contradiction to the assumption. Thus we may assume that $C_{1}=x_{1}e_{1}x_{2}e_{2}\cdots x_{3m}e_{3m}$ contains an edge in $H_{e}\setminus H$. Recall that
$H_{e}\setminus H=\{\{w,u_{1}\},\{w,u_{2}\},\ldots,\{w,u_{k}\},\{u_{1},v_{1}\},\{u_{2},v_{2}\},\ldots,\{u_{k},v_{k}\}\}.$
Thus if $C_{1}$ has an edge in $H_{e}\setminus H$, then there are exactly $4$ consecutive edges in $C$ contained in $H_{e}\setminus H$. Without loss of generality, assume $e_{1}=\{u_{1},v_{1}\}$, $e_{2}=\{w,u_{1}\}$, $e_{3}=\{w,u_{2}\}$, $e_{4}=\{u_{2},v_{2}\}$ and $x_{1}=v_{1}$, $x_{2}=u_{1}$, $x_{3}=w$, $x_{4}=u_{2}$, $x_{5}=v_{2}$. Note that all other edges $e_{5},\ldots,e_{3m}$ are edges of $H$. Then $C^{\prime}_{1}=v_{1}ev_{2}e_{5}x_{6}e_{6}\cdots x_{3m}e_{3m}$ is a Berge cycle of length $33m-$ 3 in $H$.
Now, for each $i\in\{2,3,\ldots,k\}$, the vertex $w$ does not appear in $C_{i}$ since $C_{i}$ is disjoint with $C_{1}$. Then the above argument implies that $C_{i}$ cannot contain an edge in $H_{e}\setminus H$. Thus we can easily check that $C^{\prime}_{1},C_{2},\ldots,C_{k}$ are pairwise disjoint ternary Berge cycles in $H$, implying $t(H)\geq k$ which is again a contradiction. ∎
When $k=0$, the above lemma states the following:
Corollary 5.3.
Let $H$ be a hypergraph and $e$ be an edge of $H$. If $H$ has no ternary Berge cycle, then neither does $H_{e}$.
Now we are ready to prove Theorem 1.3. This can be done by converting the original hypergraph without a ternary Berge cycle to a graph without a cycle of length 00 mod $3$, preserving the total Betti number. Then it follows from the result in [16]:
Theorem 5.4.
[16] If a graph $G$ has no induced cycle of length 00 mod $3$, then $\beta(I(G))\leq 1$.
Theorem 1.3.
If a hypergraph $H$ has no ternary Berge cycle, then $\beta(I(H))\leq 1$.
Proof.
Let $H$ be a hypergraph with no ternary Berge cycle. We claim that there is a graph $G$ with no ternary cycle with $\beta(I(G))=\beta(I(H))$. Then by Theorem 5.4, it implies that $\beta(I(H))=\beta(I(G))\leq 1$.
We apply induction on the number of edges of size at least $3$. As mentioned in Section 2, we assume that $H$ has no edge of size $1$. Thus the base case of the induction is when all edge of $H$ has the size $2$. In this case, the statement follows from Theorem 5.4. Now suppose $H$ has an edge $e$ of size at least $3$. We observe the following about $H_{e}$:
  • By the definition, the number of edges of size at least $3$ in $H_{e}$ is one less than that of $H$.
  • $\beta(I(H))=\beta(I(H_{e}))$ by Lemma 5.1 and Prop 2.1.
  • Corollary 5.3 implies that $H_{e}$ also has no ternary Berge cycle.
By the induction hypothesis, $\beta(I(H_{e}))\leq 1$, thus it follows that $\beta(I(H))\leq 1$. ∎
A natural question that arises from Theorem 1.3 is which homotopy types that $I(H)$ can have for such hypergraphs. We suspect that, if $H$ contains no ternary Berge cycles, then $I(H)$ should be either contractible or homotopy equivalent to a sphere, as in the case of graphs [13]. However, this does not immediately follow from our argument because of the existence of a complex whose suspension is homotopy equivalent to a sphere while the complex itself is not. By the double suspension theorem [5], the double suspension $\Sigma^{2}X$ of a homology $d$-sphere $X$ is homotopy equivalent to $\mathbb{S}^{d+2}$. For instance, if we consider the Poincaré homology $3$-sphere, then its fundamental group has order $120$ and hence it is not homotopy equivalent to a sphere, but its double suspension is homotopy equivalent to $\mathbb{S}^{5}$. As a conclusion, the following question still remains:
Question 5.5.
If a hypegraph $H$ has no ternary Berge cycle, then is $I(H)$ either contractible or homotopy equivalent to a sphere?
Remark 5.6.
In the proof of Theorem 1.3, one of the key part is Corollary 5.3, which says that if $H$ has no ternary Berge cycle, then neither does $H_{e}$ for any edge $e$ of $H$. It might be wondered whether the converse direction holds. That is, if $H$ has no ternary Berge cycle, then does $H_{v}$ also have no ternary Berge cycle for any vertex $v$ of $H$? Unfortunately, the answer is negative in general, and this is the reason why we consider $H_{e}$ instead of $H_{v}$ to prove Theorem 1.3. See Figure 5 for an example.
Figure 5. Although $H$ has no ternary Berge cycle, $H_{v}$ contains a triangle.
5.2. Bounded number of ternary Berge cycles
Now we investigate more general case when a hypergraph $H$ contains at most $k$ disjoint ternary Berge cycles, i.e. $t(H)\leq k$. We prove that $\beta(I(H))$ is bounded by a function of $k$. For this, we use the following Erdős–Pósa theorem [15] for cycles of length 00 mod $m$.
Theorem 5.7.
[15] There is a function $g_{m}:\mathbb{N}\to\mathbb{N}$ satisfying the following. If a graph $G$ contains at most $k$ vertex-disjoint cycles of length 00 mod $m$, then there are at most $g_{m}(k)$ vertices hitting all cycles of length 00 mod $m$.
The current best known bound for $g_{m}(k)$ is $g_{m}(k)=O_{m}(k\log k)$ by [2].
Theorem 5.8.
For all graph $G$ with $t(G)\leq k$, we have $\beta(I(G))\leq 2^{g_{3}(k)}=2^{O(k\log k)}$.
Proof.
For every vertex $v$ of $G$, we have
$I(G)=I(G-v)\cup I(G-N_{G}(v))\text{ and }I(G-v)\cap I(G-N_{G}(v))=I(G-N_{G}[v]).$
Since $v$ is an isolated vertex of $G-N_{G}(v)$, we note that $I(G-N_{G}(v))$ is a cone with apex $v$, thus it is contractible. This implies $\tilde{\beta}_{i}(I(G-N_{G}(v)))=0$ for all $i$. Then by Lemma 2.3, we obtain
$\tilde{\beta}_{i}(I(G))\leq\tilde{\beta}_{i}(I(G-v))+\tilde{\beta}_{i-1}(I(G-N_{G}[v]))$
for all $i$. Thus we have
$\beta(I(G))\leq\beta(I(G-v))+\beta(I(G-N_{G}[v])$
for any vertex $v$.
Now, for disjoint vertex subsets $X$ and $Y$, let $G(X|Y)$ be the graph obtained from $G$ by deleting all vertices in $N_{G}[X]\cup Y$, that is, $G(X|Y)=G-N_{G}[X]-Y$. For convenience, we omit curly brackets in the notation $G(X|Y)$, for instance we write $G(v|Y)$ instead of $G(\{v\}|Y)$. Then the above inequality can be written as
(2) $\beta(I(G))\leq\beta(I(G(v|\varnothing)))+\beta(I(G(\varnothing|v)))$
for any vertex $v$.
Given a vertex subset $U=\{u_{1},u_{2},\ldots,u_{m}\}$, by the repeated usage of (2), we obtain
$\beta(I(G))$ $\leq\beta(I(G(u_{1}|\varnothing)))+\beta(I(G(\varnothing|u_{1})))$
$\leq\beta(I(G(u_{1},u_{2}|\varnothing))+\beta(I(G(u_{1}|u_{2})))+\beta(I(G(u_{2}|u_{1}))+\beta(I(G(\varnothing|u_{1},u_{2})))$
$\leq\cdots\leq\sum_{X\subset U}\beta(I(G(X|U\setminus X))).$
By Theorem 5.7, we can take a vertex subset $W$ with $|W|\leq g_{3}(k)$ hitting all ternary cycles of $G$. Then we have
$\beta(I(G))\leq\sum_{X\subset W}\beta(I(G(X|W\setminus X))).$
On the other hand, since $N_{G}[X]\cup(W\setminus X)\supset W$, the graph $G(X|W\setminus X)=G-N_{G}[X]-(W\setminus X)$ is a ternary graph. Then by Theorem 5.4, we have $\beta(I(G(X|W\setminus X)))\leq 1$ for any $X\subset W$, and hence we obtain
$\beta(I(G))\leq\sum_{X\subset W}\beta(I(G(X|W\setminus X)))\leq 2^{|W|}=2^{g_{3}(k)}=2^{O(k\log k)}$
as desired. ∎
By applying Lemma 5.1, Lemma 5.2 and Proposition 2.1, we can obtain the following hypergraph version as a consequence of Theorem 5.8.
Corollary 5.9.
For all hypergraph $H$ with $t(H)\leq k$, we have $\beta(I(H))\leq 2^{O(k\log k)}$.
Remark 5.10.
A trivial lower bound for Theorem 5.8 (also for Corollary 5.9) is $2^{k}$ because we have $\beta(I(G))=2^{k}$ when $G$ is a disjoint union of $k$ ternary cycles.
In general, if $G=G_{1}\dot{\sqcup}G_{2}$, i.e. $G$ is a disjoint union of two graphs $G_{1},G_{2}$, then it is known that $I(G)=I(G_{1})*I(G_{2})$ where the join of two simplicial complexes $X,Y$ are defined as follows:
$X*Y=\{\sigma\cup\tau\mid\sigma\in X,\tau\in Y\}.$
Furthermore, by the Künneth formula, it is also known that
$\tilde{\beta}_{r+1}(X*Y)=\sum_{i+j=r}\tilde{\beta}_{i}(X)\tilde{\beta}_{j}(Y)$
for all $r$, and it implies $\beta(X*Y)=\beta(X)\beta(Y)$. Hence we obtain $\beta(I(G))=\beta(I(G_{1}))\beta(I(G_{2}))$ when $G=G_{1}\dot{\sqcup}G_{2}$.
Thus if $G=\dot{\bigsqcup}_{i=1}^{k}C_{i}$ where each $C_{i}$ is a ternary cycle, then we obtain
$\beta(I(G))=\beta(I(C_{1}))\beta(I(C_{2}))\cdots\beta(I(C_{k}))=2^{k}$
since we know that $\beta(I(C))=2$ for every ternary cycle $C$.
Acknowledgement
The author thanks Sang-il Oum for his help in finding the current best known bound on the Erdős–Pósa theorem for cycles of length 00 mod $m$. The author also thank Minki Kim for helpful s and discussions throughout the research of this paper. This study was financially supported by Chonnam National University (Grant number: 2024-0467). The author was also supported by the Institute for Basic Science (IBS-R029-C1).